What is the slope of the line passing through (0,-1), and (3,-2)?
answer in the book is 3 but my answer is -1/3. Please tell why my answer is different from the book
Thank you

Answer by richwmiller(17219)

(y1-y2)/x1-x2)=m
-1-(-2)/0-3=
-1+2/-3=
1/-3=
-1/3
Now doing it from y2
-2-(-1)/3-0
-2+1/3
-1/3
It should not matter whether you subtract y1 from y2 or y2 from y1 as long as you do the same with x. We are measuring the difference which is the change which is the same whether you measure it up or down.

The book is wrong. You are right
You will have fun if the teacher wasn't aware of error in the book.
It very common for the books to have wrong answers.
First is a problem with typos, just plain typing mistakes.
Those are actually not too common.
What usually happens is that there are last minute changes to the problem for what ever reason and they forget to change the answer key.
Sometimes they completely replace the problem or change some numbers.
Your teacher should be aware of this continual nuisance.
You might get extra credit for sticking to your guns.
If I were the teacher I would give extra credit to any one who could find mistakes in the answers.
It means you are paying attention.